# Domain Generalization: A Tale of Two ERMs

### v1

@@ Appendix A Additional literature review
Empirically, Gulrajani and Lopez-Paz (2021) first emphasized that a well-tuned empirical risk minimization (ERM) baseline outperforms many domain generalization (DG) methods on vision benchmarks. Similar patterns were later observed on the WILDS benchmark (Koh et al., 2021), and again in the context of federated domain generalization by Bai et al. (2024).
On the theoretical front, Rosenfeld et al. (2021) and Gouk et al. (2024) studied function classes of the form $f:\mathcal{X}\to\mathcal{Y}$ and concluded that, under common assumptions, ERM cannot be fundamentally outperformed (e.g., in terms of minimax risk). The recent work by Dwork et al. (2025) analyzed the optimality of the “min-max” ERM with another performance measure of interest.
@@
While most DG methods restrict themselves to using only the input $x$ at inference time, some methods explore the use of unlabeled test-domain data. Several DG methods attempt to exploit unlabeled test data to improve generalization (Blanchard et al., 2011; Muandet et al., 2013; Zhang et al., 2021). A closely related setting is unsupervised domain adaptation (UDA), where unlabeled test data are used to adapt models to the test domain. Unlike DG, UDA assumes access to target-domain data at training time and typically requires learning a separate model per test domain (Sun and Saenko, 2016; Ganin et al., 2016).
Although promising in principle, the practical benefits of using unlabeled data remain mixed. A large-scale study by Sagawa et al. (2022) evaluating methods across ten diverse datasets found that incorporating unlabeled data frequently failed to improve upon strong ERM baselines. These findings reinforce the need for a more precise understanding of when and how unlabeled data can contribute to domain generalization.
@@
Appendix B A Simple Example of Covariate Shift
The example below show that even if there is no decision-theoretic gain of DI-ERM under covariate shift (Corollary˜1), it may still have practical advantage when considering learning over a restricted function class.
Figure 2: Illustration of Example 1, where $R^{*}_{\mathrm{pool},\,\mathcal{G}}>R^{*}_{\mathrm{DG},\,\mathcal{F}}$.
Example 1 (Covariate shift without posterior drift)
@@ Appendix B A Simple Example of Covariate Shift
$f^{*}_{\mathrm{DG}}(x,m)$ $=\operatorname{sign}(x-4m+3)\quad\Longrightarrow f^{*}_{\mathrm{pool}}=f^{*}_{\mathrm{DG}}$
therefore $R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}=0$. The model classes are linear classifiers
@@
$\mathcal{F}$ can realize $f^{*}_{\mathrm{DG}}$ with a bias term that depends on $m$, giving $R^{*}_{\mathrm{DG},\,\mathcal{F}}=0$. However, a predictor in $\mathcal{G}$ can only choose a single threshold, and the optimal one is
@@
Therefore, $R^{*}_{\mathrm{pool},\,\mathcal{G}}>R^{*}_{\mathrm{DG},\,\mathcal{F}}$, even though $R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}$.
This toy construction mirrors image classification task across different styles: each style (domain) has a separate support, so the Bayes classifier is the same with or without $m$, yet $m$ still helps within a restricted model class. This is experimentally illustrated in Section˜6 .
@@ Appendix C Proofs
C.1 Proof of Proposition˜1
Proposition (Risk Hierarchy)  $R^{*}_{\mathrm{pool}}\geq R^{*}_{\mathrm{DG}}\geq R^{*}_{\mathrm{full}}.$
Proof. Although this result follows from Theorems 1 and 2, it can be proved more directly as follows:
@@ C.1 Proof of Proposition˜1
$\geq\inf_{f:\mathcal{X}\times\mathcal{M}\to\mathcal{Y}}\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(X,M)\neq Y}\right]=R^{*}_{\mathrm{DG}}$
@@
To see the final step, note that the $f(x,m,d)$ achieving the minimum in the next-to-last line is $\mathop{\mathrm{argmax}}_{k}{\mathbb{P}}(Y=k\mid X=x,M=m,D=d)$. By Assumption (1), this is equal to $\mathop{\mathrm{argmax}}_{k}{\mathbb{P}}(Y=k\mid X=x,D=d)$, which achieves the minimum in the last line.
C.2 Proof of Theorem˜1
@@
$\mathbb{E}_{X,M}\left[\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]\leq R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}\leq\mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
@@
$f^{*}_{\mathrm{pool}}(X)=f^{*}_{\mathrm{DG}}(X,M)\quad\text{ almost surely w.r.t. }P_{XM}\implies R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}.$
Proof. The gap in the two risks can be expressed as
$R^{*}_{\mathrm{pool}}$ $-R^{*}_{\mathrm{DG}}$
$=$ $\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)|X,M\right)\right]\right]-\mathbb{E}_{X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X\right)\right]$
$=$ $\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)|X,M\right)\right]\right]-\mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X,M\right)\right]\right]$
$=$ $\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)|X,M\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X,M\right)\right]\right]$
Notice that for any $x,m$, if $f^{*}_{\mathrm{DG}}(x,m)=f^{*}_{\mathrm{pool}}(x)$, then the pointwise difference of the conditional probabilities inside the expectation above must be zero.
@@
$\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(X,M)|X,M\right)$ $=\max_{k}P(Y=k|X,M),$
@@
$\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X,M\right)$ $\leq\operatorname{2nd}\max_{k}P(Y=k|X,M).$
It thus follows that.
$\gamma(x,m)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,m)}$ $\leq\mathbb{P}\left(Y=f^{*}_{\mathrm{DG}}(x,m)|X=x,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(x)|X=x,M=m\right)$
$\leq\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,m)}.$
The inequalities in the theorem statement now follow.
From the lower and upper bound above, we can directly get the sufficient condition that
$f^{*}_{\mathrm{pool}}(X)=f^{*}_{\mathrm{DG}}(X,M)\quad\text{ almost surely w.r.t. }P_{XM}\implies R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DG}}.$
C.3 Proof of Proposition˜2
@@
$\inf_{P_{XYMD}\in\Pi(\gamma,\epsilon)}\Big[R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}\Big]\geq\frac{\gamma\cdot\epsilon}{2}$
Proof. From the lower bound in Theorem˜1, we have
@@
$\geq\gamma\ \mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$ $\because\text{margin assumption in }\Pi(\gamma,\epsilon)$
$=\gamma\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]\right]$
@@
$\forall x,\quad\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,M)}\right]\geq\frac{1}{2}\ \mathbb{E}_{M,M^{\prime}|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(x,M)\neq f^{*}_{\mathrm{DG}}(x,M^{\prime})}\right],$
@@
$\pi_{k}(x)=\mathbb{P}\left(f^{*}_{\mathrm{DG}}(x,M)=k|X=x\right),$
@@
$\mathbb{E}_{M,M^{\prime}|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(x,M)\neq f^{*}_{\mathrm{DG}}(x,M^{\prime})}\right]$ $=\mathbb{P}\left(f^{*}_{\mathrm{DG}}(x,M)\neq f^{*}_{\mathrm{DG}}(x,M^{\prime})|X=x\right)$
$=\sum_{k}\mathbb{P}\left(f^{*}_{\mathrm{DG}}(x,M)=k,f^{*}_{\mathrm{DG}}(x,M^{\prime})\neq k|X=x\right)$
@@
$\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,M)}\right]$ $=\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(x,M)\neq k_{0}}\right]$
@@
$\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(x,M)}\right]\geq\frac{1}{2}\ \mathbb{E}_{M,M^{\prime}|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(x,M)\neq f^{*}_{\mathrm{DG}}(x,M^{\prime})}\right].$
@@
$\mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]\right]$ $\geq\frac{1}{2}\mathbb{E}_{X}\left[\mathbb{E}_{M,M^{\prime}|X}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(X,M)\neq f^{*}_{\mathrm{DG}}(X,M^{\prime})}\right]\right]$
$=\frac{1}{2}P_{X,M,M^{\prime}}\Big(f^{*}_{\mathrm{DG}}\big(X,M\big)\neq f^{*}_{\mathrm{DG}}\big(X,M^{\prime}\big)\Big)$

### v2

@@ Appendix A Additional literature review
Empirically, Gulrajani and Lopez-Paz [2021] first emphasized that a well-tuned empirical risk minimization (ERM) baseline outperforms many domain generalization (DG) methods on vision benchmarks. Similar patterns were later observed on the WILDS benchmark [Koh et al., 2021], and again in the context of federated domain generalization by Bai et al. [2024].
On the theoretical front, Rosenfeld et al. [2021] and Gouk et al. [2024] studied function classes of the form $f:\mathcal{X}\to\mathcal{Y}$ and concluded that, under common assumptions, ERM cannot be fundamentally outperformed (e.g., in terms of minimax risk). The recent work by Dwork et al. [2025] analyzed the optimality of the “min-max” ERM with another performance measure of interest.
Invariant feature learning.
A large body of work in domain generalization aims to learn features that are invariant, or transferable, across domains. CORAL [Sun and Saenko, 2016] aligns second-order statistics of feature representations across domains. DANN [Ganin et al., 2016] employs adversarial training to encourage domain-invariant features. Invariant Risk Minimization (IRM) [Arjovsky et al., 2019] seeks a representation for which a single classifier is simultaneously optimal across all training domains. Bui et al. [2021] disentangle domain-invariant and domain-specific features and discard the latter at inference time. For more comprehensive overviews, we refer the reader to the surveys by Wang et al. [2022a] and Zhou et al. [2023].
@@
While most DG methods restrict themselves to using only the input $x$ at inference time, some methods explore the use of unlabeled test-domain data. Several DG methods attempt to exploit unlabeled test data to improve generalization [Blanchard et al., 2011, Muandet et al., 2013, Zhang et al., 2021]. A closely related setting is unsupervised domain adaptation (UDA), where unlabeled test data are used to adapt models to the test domain. Unlike DG, UDA assumes access to target-domain data at training time and typically requires learning a separate model per test domain [Sun and Saenko, 2016, Ganin et al., 2016].
Although promising in principle, the practical benefits of using unlabeled data remain mixed. A large-scale study by Sagawa et al. [2022] evaluating methods across ten diverse datasets found that incorporating unlabeled data frequently failed to improve upon strong ERM baselines. These findings reinforce the need for a more precise understanding of when and how unlabeled data can contribute to domain generalization.
@@
Learning Using Privileged Information (LUPI)
Our formulation of DG also is related to the learning using privileged information (LUPI) framework [Vapnik and Vashist, 2009, Sharmanska et al., 2014, Vapnik et al., 2015], where additional side information is available during training. Recent works have explored LUPI in domain adaptation settings [Breitholtz et al., 2024].
Two key differences distinguish our setting from LUPI:
  • LUPI typically provides different privileged information for each sample, whereas in our DG setup, samples from the same domain share the same domain metadata.
  • Crucially, we assume access to domain metadata at test time, while LUPI typically assumes side information is only available during training. This difference is fundamental to our theoretical contributions: LUPI guarantees smaller sample complexity, while our setting enables strictly smaller Bayes risk.
Appendix B A Simple Example of UBC
The example below show that even if there is no decision-theoretic gain of DI-ERM under UBC (Corollary 1), it may still have practical advantage when considering learning over a restricted function class.
Figure 2: Illustration of Example 1, where $R^{*}_{\mathrm{pool},\,\mathcal{G}}>R^{*}_{\mathrm{DI},\,\mathcal{F}}$.
Example 1.
@@ Appendix B A Simple Example of UBC
$f^{*}_{\mathrm{DI}}(x,m)$ $=\operatorname{sign}(x-4m+3)\quad\Longrightarrow f^{*}_{\mathrm{pool}}=f^{*}_{\mathrm{DI}}$
therefore $R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DI}}=0$. The model classes are linear classifiers
@@
$\mathcal{F}$ can realize $f^{*}_{\mathrm{DI}}$ with a bias term that depends on $m$, giving $R^{*}_{\mathrm{DI},\,\mathcal{F}}=0$. However, a predictor in $\mathcal{G}$ can only choose a single threshold, and the optimal one is
@@
Therefore, $R^{*}_{\mathrm{pool},\,\mathcal{G}}>R^{*}_{\mathrm{DI},\,\mathcal{F}}$, even though $R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DI}}$.
This toy construction mirrors image classification task across different styles: each style (domain) has a separate support, so the Bayes classifier is the same with or without $m$, yet $m$ still helps within a restricted model class. This is experimentally illustrated in Section 6 .
@@ Appendix C Proofs
C.1 Proof of Proposition 1
Proposition (Risk Hierarchy)  $R^{*}_{\mathrm{pool}}\geq R^{*}_{\mathrm{DI}}\geq R^{*}_{\mathrm{full}}.$
Proof.
Although this result follows from Theorems 1 and 3, it can be proved more directly as follows:
@@ C.1 Proof of Proposition 1
$\geq\inf_{f:\mathcal{X}\times\mathcal{M}\to\mathcal{Y}}\mathbb{E}_{X,Y,M,D}\left[\mathbbm{1}_{f(X,M)\neq Y}\right]=R^{*}_{\mathrm{DI}}$
@@
To see the final step, note that the $f(x,m,d)$ achieving the minimum in the next-to-last line is $\mathop{\mathrm{argmax}}_{k}{\mathbb{P}}(Y=k\mid X=x,M=m,D=d)$. By Assumption (1), this is equal to $\mathop{\mathrm{argmax}}_{k}{\mathbb{P}}(Y=k\mid X=x,D=d)$, which achieves the minimum in the last line. ∎
C.2 Proof of Theorem 1
@@
$\mathbb{E}_{X,M}\left[\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]\leq R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DI}}\leq\mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]$
@@
$f^{*}_{\mathrm{pool}}(X)=f^{*}_{\mathrm{DI}}(X,M)\quad\text{ almost surely w.r.t. }P_{XM}\implies R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DI}}.$
Proof.
The gap in the two risks can be expressed as
$R^{*}_{\mathrm{pool}}$ $-R^{*}_{\mathrm{DI}}$
$=$ $\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(X,M)|X,M\right)\right]\right]-\mathbb{E}_{X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X\right)\right]$
$=$ $\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(X,M)|X,M\right)\right]\right]-\mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X,M\right)\right]\right]$
$=$ $\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(X,M)|X,M\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X,M\right)\right]\right]$
Next, we analyze the expression in a pointwise manner (aka, for every $x$ and $m$).
Notice that for any $x,m$, if $f^{*}_{\mathrm{DI}}(x,m)=f^{*}_{\mathrm{pool}}(x)$, then the pointwise difference of the conditional probabilities inside the expectation above must be zero.
@@
$\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(X,M)|X,M\right)$ $=\max_{k}\mathbb{P}\left(Y=k|X,M\right),$
@@
$\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(X)|X,M\right)$ $\leq\operatorname{2nd}\max_{k}\mathbb{P}\left(Y=k|X,M\right).$
Thus concluding the two scenarios, the lower bound follows that
$\gamma(x,m)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DI}}(x,m)}$ $\leq\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(x,m)|X=x,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(x)|X=x,M=m\right).$
The upper bound can be bounded by the difference of the two optimal classifiers
$\mathbb{P}\left(Y=f^{*}_{\mathrm{DI}}(x,m)|X=x,M=m\right)-\mathbb{P}\left(Y=f^{*}_{\mathrm{pool}}(x)|X=x,M=m\right)$ $\leq\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DI}}(x,m)}.$
Taking the expectation over $X$ and $M$, we get the upper and lower bounds in the theorem statement
$\mathbb{E}_{X,M}\left[\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]\leq R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DI}}\leq\mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]$
From the (pointwise) lower and upper bound above, we can get the sufficient condition that
$f^{*}_{\mathrm{pool}}(X)=f^{*}_{\mathrm{DI}}(X,M)\quad\text{ almost surely w.r.t. }P_{XM}\implies R^{*}_{\mathrm{pool}}=R^{*}_{\mathrm{DI}}.$
C.3 Proof of Theorem 2
@@
$\inf_{P_{XYMD}\in\Pi(\gamma,\epsilon)}\Big[R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DI}}\Big]\geq\frac{\gamma\cdot\epsilon}{2}$
Proof.
From the lower bound in Theorem 1, we have
@@
$\geq\gamma\ \mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]$ $\because\text{margin assumption in }\Pi(\gamma,\epsilon)$
$=\gamma\ \mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]\right]$
@@
$\forall x,\quad\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DI}}(x,M)}\right]\geq\frac{1}{2}\ \mathbb{E}_{M,M^{\prime}|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DI}}(x,M)\neq f^{*}_{\mathrm{DI}}(x,M^{\prime})}\right],$
@@
$\pi_{k}(x)=\mathbb{P}\left(f^{*}_{\mathrm{DI}}(x,M)=k|X=x\right),$
@@
$\mathbb{E}_{M,M^{\prime}|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DI}}(x,M)\neq f^{*}_{\mathrm{DI}}(x,M^{\prime})}\right]$ $=\mathbb{P}\left(f^{*}_{\mathrm{DI}}(x,M)\neq f^{*}_{\mathrm{DI}}(x,M^{\prime})|X=x\right)$
$=\sum_{k}\mathbb{P}\left(f^{*}_{\mathrm{DI}}(x,M)=k,f^{*}_{\mathrm{DI}}(x,M^{\prime})\neq k|X=x\right)$
@@
$\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DI}}(x,M)}\right]$ $=\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DI}}(x,M)\neq k_{0}}\right]$
@@
$\mathbb{E}_{M|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DI}}(x,M)}\right]\geq\frac{1}{2}\ \mathbb{E}_{M,M^{\prime}|X=x}\left[\mathbbm{1}_{f^{*}_{\mathrm{DI}}(x,M)\neq f^{*}_{\mathrm{DI}}(x,M^{\prime})}\right].$
@@
$\mathbb{E}_{X}\left[\mathbb{E}_{M|X}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(x)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]\right]$ $\geq\frac{1}{2}\mathbb{E}_{X}\left[\mathbb{E}_{M,M^{\prime}|X}\left[\mathbbm{1}_{f^{*}_{\mathrm{DI}}(X,M)\neq f^{*}_{\mathrm{DI}}(X,M^{\prime})}\right]\right]$
$=\frac{1}{2}P_{X,M,M^{\prime}}\Big(f^{*}_{\mathrm{DI}}\big(X,M\big)\neq f^{*}_{\mathrm{DI}}\big(X,M^{\prime}\big)\Big)$